If a rectangle's length is 7 more than 4 times the width and the perimeter is 54 what are the dimensions of the rectangle?
Answer by josgarithmetic(39617)

Exactly the same as question number 884637.

L=7+4w and 2w+2L=54.

You can continue through the steps starting with 2w+2(7+4w)=54.